A uniform slab of dimensional 10 cm x 10 cm x 1 cm  is kept between two heat reservoirs at temperatures  10°C and 90°C. The larger surface areas touch the reservoirs . The thermal conductivity  of the material is 0.80 W/m°C. Find the amount of heat flowing through the slab per min find the amount of heat flowing through the slap per minute 

Grade:12

1 Answers

Eshan
5 years ago
\dfrac{dH}{dt}=K\dfrac{A\Delta T}{t}
=0.8\times \dfrac{(0.1\times 0.1)(90-10)}{0.01}W
=64W=3.84kJ/min